# Interactions across multiple games: cooperation, corruption, and organizational design

### v1

@@ 1 Introduction
We study when and how the structure and design of organizations can encourage socially beneficial forms of cooperation while inhibiting harmful forms of cooperation. We model teams that face a random arrival of games over time. When a game arrives, each team member chooses whether to cooperate with others. In any given game, a team member prefers overall team cooperation to a lack of cooperation but individually prefers not to cooperate. Thus, the games have a prisoners’ dilemma-like structure. This is not a simple repeated game: the stage games take two different forms. In some of the games, cooperation is beneficial to society (“good games”); in other games, cooperation is harmful (“bad games”). Socially optimal equilibria are those in which team members cooperate in good games only. Yet as Tirole (1988) conjectured, “…it may be worth tolerating some minor acts of detrimental co-operation to allow trust between employees to develop.” We show that the team-design problem is nontrivial because of interdependencies between the different stage games: equilibrium structure depends on the arrival frequencies of both good and bad games as well as their payoff structures.
A key design dimension is team durability. Some organizations, such as the military or police, keep teams together over time to foster cooperation. Others—e.g., diplomacy, retail, logistics, and services—frequently reshuffle members across teams, eliminating the repeated interaction necessary for cooperation. In which settings should teams be kept together versus reshuffled? Again, this issue is nontrivial as Tirole (1988) anticipated: “Fighting the detrimental cooperation by keeping the employees’ relationship short may also kill the beneficial co-operation.” We formalize this insight and identify conditions that make detrimental cooperation necessary to sustain socially beneficial cooperation.
@@
We also consider dynamic versions of the assignment of teams to tasks. Teams may be completing a certain task for some duration of time before being moved to another; they may also be moved in response to, say, an opportunity to cooperate in a situation where it is socially beneficial, e.g., after police officers find themselves in a dangerous spot. Dynamically and reactively reassigning teams to tasks further allows the organization to affect the random stream of opportunities to cooperate. We fully characterize the optimal structure of such assignments for a special case—when it is sufficiently easy to sustain cooperation in the bad game.
@@
Game theory, team theory, and mechanism design generally take the composition of teams and the set of tasks that they face as exogenous. [^3] There is a literature that examines how the diversity of teams affects their performance potential (e.g., see Page (2008)), but that is based on team heterogeneity, which is not considered here, and that literature does not consider any of the incentive issues or dynamics considered here. To our knowledge, this is the first paper to analyze the interaction of incentives across games such that cooperation is socially beneficial in some but not others. Importantly, this provides a first analysis of treating the assignments of players to teams, and teams to games, as endogenous. This provides new angles for understanding team and organizational design.
@@
Our model is a stochastic game (introduced by Shapley, 1953; see Solan and Vieille, 2015 for a short review). Our model’s special structure—the players’ current actions do not affect the distribution over subsequent games—may be useful for extensions. This environment allows us to characterize equilibria, which are typically difficult to solve for (Solan and Vieille, 2015).
@@ 2 The Model
Player 1 C $(c,c)$ $(-a,c+d)$
N $(c+d,-a)$ $(0,0)$
where $c,d>0$ and $a>c+d$. All players cooperating maximizes the total payoff, but a player can benefit by unilaterally deviating. The condition that $a>c+d$ ensures that the game satisfies Properties 5 (since then $0>c+d-a$) and 6 (since then $c+d-c\leq 0-(-a)$). The resulting cooperation benefit and deviation temptation are $c_{g}\equiv c$ and $d_{g}\equiv d$, respectively.
@@ 2.1 Cooperation
The different kinds of cooperation can be partially ordered by some social criterion. Let $V_{G}>0$ denote the social benefit from cooperation in game $G$, the good game, from the perspective of a social planner or designer. For example, this could be the social benefits to a society’s broader population of having a military or police force that cooperates when in dangerous situations, or the profits to a company from having a team cooperate on a productive task, etc. Similarly, $V_{B}<0$ denotes the social harm from cooperation in the bad game $B$. Thus, cooperation only in game $G$ is clearly most preferred; we call this optimal cooperation. Cooperation only in game $B$ is clearly the worst. The social value of total cooperation, $p_{G}V_{G}+p_{B}V_{B}$, and the social value of no cooperation, 00, lie in between the best and worst outcomes; their ordering depends on the relative social impact of cooperation in the two games and on their probabilities.
@@ 2.2 Player-Optimal Equilibria
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{B}/c_{B}$$\ x$region 2region 1region 3
@@
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{G}/c_{G}$$\ x$region 2region 1region 4
@@ 3 How Cooperation Depends on Team Durability
We model the reshuffling of teams (reducing team durability) as a probability $r$ with which the stochastic game of a team ends after each period. Given the continuum of workers, we assume that reshuffled players never face a past teammate and that there is no spreading of information and that players only condition on histories with their current team. [^7] This greatly simplifies the analysis; see, e.g., Ali and Miller (2016b) for a treatment where such contagion is possible. Our simplification also requires sufficiently many workers so that completely new teams can be formed and anonymity of workers outside their teams (deviations can be punished only by current team members). These conditions are more likely to hold—at least approximately—in large organizations, potentially leading to differences by organization size that we leave for future research. Consequently, a reshuffling rate $r$ is equivalent to changing the discount factor $\delta$ to $(1-r)\delta$; essentially, the designer can induce any effective discount factor of at most $\delta$ by varying team durability. Thus, formally, we characterize optimal reshuffling by examining the player optimal equilibria for a given team (with histories only applying to that team) as a function of discount factor $(1-r)\delta$.
@@
When any of the above conditions fail, there is a trade-off between two evils: accept corruption in bad games or face cooperation in neither game. If total cooperation is feasible then it is preferred to no cooperation if and only if $p_{G}V_{G}+p_{B}V_{B}>0$. In policing, for example, productive team cooperation is vital to society—$V_{G}$ is high. However, because police misconduct imposes significant harm, $V_{B}$ can also be large (Mollen Commission, 1994). Thus, whether frequent rotations or long-term partnerships are optimal depends on the relative frequency of situations in which (a) police need to back each other up or perform other important team duties versus (b) they can be corrupt or abusive.
@@ 4 Assigning Teams to Tasks
A team’s history at time $\tau$ is now augmented with the history of assigned tasks, i.e., it is a vector $[(t_{\tau^{\prime}},g_{\tau^{\prime}},\mathbf{a}_{\tau^{\prime}})]^{\tau-1}_{\tau^{\prime}=0}$ where $t_{\tau^{\prime}}$ is the task faced at time $\tau^{\prime}$. Let $\mathcal{H}_{\tau}$ denote all such histories and $\mathcal{H}\equiv\cup_{\tau}\mathcal{H}_{\tau}$. A player’s strategy indicates what the player who falls under a particular $\nu$ does as a function of the history, i.e., $\sigma:\mathcal{H}\times\{G,B\}\rightarrow\Delta(C,N)$, similar to before. Again, we track histories only within a team and consider equilibria that depend only on current team histories.
@@ 5.2.3 The Continuous Time Limit
We conclude our discussion on designer-optimal reactive team structures by considering how they depend on the length of a time period. In particular, we consider the continuous limit of our model; i.e., when time periods become infinitesimally small and, with them, the probability of a game occurring and the potential frequency of reassignments.
@@
We first note that as $T$ decreases, the designer’s payoff from an optimal static team structure increases. The reason is that teams do not need to wait for at least another period of length $1$ for another game after a game arrives; instead, the next game can arrive after a time interval of time $T<1$. Given any static assignment, the incentives to cooperate increase as $T$ decreases. Hence, the share of bad tasks can be reduced, and the social value improves.
Now consider reactive team structures and the case where the designer observes good games occurring (Section 5.2.1). The optimal reactive team structure does not change substantively: Consider a reactive assignment in which players always cooperate, which is part of an optimal reactive team structure. In the limit as $T\rightarrow 0$, following a good game, teams spend some fixed amount of time on the bad task before returning to the good task, where they wait until a good game arrives, as before. The only difference from before is that there is less of a “waiting period” after the arrival of a game as $T$ goes to 0, therefore improving incentives from the faster arrival rates of games. Furthermore, as was true in the $T=1$ case, the social value is higher when reactive assignments are possible compared to static assignments, as the designer can wait for the arrival of a good game. [^18] Thus, combining this conclusion with the observation in the previous paragraph, both the reactive and static assignments are improving, but the reactive is strictly better.
@@ 6 Concluding Remarks
In closing, we discuss a few extensions. In our model, the set of tasks is fixed. However, when supporting optimal cooperation is challenging, it can be advantageous to introduce new tasks that provide additional benefits from cooperation. These new tasks might not benefit society or directly increase organizational productivity, but they might help sustain cooperation in good games. For instance, scheduling “team-building” exercises, social events, and non-production related contests in which team members must cooperate with each other can help induce cooperation in other games as well. Many organizations use such team-building to enhance the loyalty that people feel towards the units in which they are involved, fostering cooperation. Although such events may be time-consuming and thus reduce the time that teams spend directly on productive activities, they may end up enhancing productivity and substitute for bad games by enabling cooperation in good games when that might have otherwise been infeasible.
There may also be individual heterogeneity of team members in their preference to engage in collusive behavior. Adding individuals with a strong inclination to defect on bad games could curb collusion but could also undermine productive cooperation. Similarly, “changing the culture” of an organization and increasing awareness of the negative consequences of collusion—that is, shifting team members’ preferences to be more averse to collusion on bad tasks—may erode the informal incentives that sustain cooperation on good tasks and productive teamwork.
Lastly, surveillance and other technologies also play an important role and may be modeled explicitly. Improved monitoring can make cooperation directly enforceable in good games and punishable in bad ones, thereby making incentive issues less relevant. Thus, organizations can benefit from investing in technologies to monitor behavior. For example, if cooperation in the good game can be directly enforced by a monitoring technology, reshuffling teams to reduce residual collusion makes good sense. However, Section 4 shows that an organization can benefit from the games’ strategic interdependencies. For example, consider an organization that prefers total cooperation to no cooperation. If no cooperation in the bad game is directly enforced via some new technology, then cooperation in the good game may be impaired as players can now only punish deviation from cooperation via that game and not the bad game. All cooperation may thereby break down.
@@ Appendix B Proofs
Consider the full-specialization team structure $\lambda$ with $\lambda(\nu^{g}(t_{g}))=q_{g}$ for $t\in\{G,B\}$, $r_{G}=0$, and $r_{B}=1$, and first suppose $d_{G}\leq\frac{\delta}{1-\delta}a_{G}c_{G}$. Consider a team assigned to $\nu^{G}$. Property 2 implies that on-path aggregate stage-game payoffs are maximized when every player chooses grim-trigger strategy ‘Always cooperate unless any player did not cooperate in the past, at which point always do not cooperate.’ Furthermore, the resulting strategy profile constitutes an equilibrium: Facing the good game, the only game players of such team face, no player benefits from deviating since not cooperating leads to an increase in the stage-game payoff of $d_{G}$ and a decrease in future payoffs of $\frac{\delta}{1-\delta}a_{G}c_{G}$ and by assumption and since not cooperating if someone deviated in the past is a Nash equilibrium (an implication of Property 3), i.e., the punishment stage is a Nash equilibrium. Consider now a team assigned to $\nu^{B}$. As $r_{B}=1$, players’ discount factor is 00 and so they play the unique stage game equilibrium, which is no cooperation. Thus, there is cooperation only on good games.
@@
Consider a class of reactive assignments $m$ parameterized by $w_{gg^{\prime}}\in[0,1]$ for $g,g^{\prime}\in\{G,B\}$ defined as follows. Consider any state $[(t_{\tau},g_{\tau})]_{\tau=0}^{K}$ with current task $t_{K}=t_{g}$; then let $m([(t_{\tau},g_{\tau})]_{\tau=0}^{K})(t_{g}^{\prime})=w_{gg^{\prime}}$ for $w_{gg^{\prime}}$ such that $w_{gG}+w_{gB}=1$ for all $g\in\{G,B\}$. Note that for $w_{Gg}=w_{Bg}=\nu^{coop}(t_{g})$, then $m$ corresponds to the same static assignment as $\nu^{coop}$.
@@
Let $\overline{g},\underline{g}$ be such that $d_{\overline{g}}>d_{\underline{g}}$, i.e., it is harder to incentivize cooperation in game $\overline{g}$. The conditions of the proposition imply $a_{G}c_{G}<a_{B}c_{B}$. Therefore, in an assignment $m$ that is not the same static assignment as $\nu^{coop}$ the average benefit from future cooperation can be made strictly larger when teams are in game $\overline{g}$ than when they are in game $\underline{g}$ by sending them more frequently to the bad task when in $\overline{g}$, i.e., by increasing $w_{\overline{g}B}$. Furthermore, this can be done while maintaining the same unconditional benefit from future cooperation in this modified $m$ and original $\nu^{coop}$ by decreasing $w_{\underline{g}B}$ according to Equation 8. By continuity and if $d_{\overline{g}}$ is strictly less than $d_{\underline{g}}$, there are thus values of $w_{BG},w_{GB}$ (with $w_{BB}=1-w_{BG},w_{GG}=1-w_{GB}$) such that the benefit from future cooperation is strictly greater than the deviation payoff in both tasks, which allows the designer to strictly improve the social value, as in the proof of Lemma 3.

### v2

@@ 1 Introduction
We study when and how the structure and design of organizations can encourage socially beneficial forms of cooperation while inhibiting harmful forms of cooperation. We model teams that face a random arrival of games over time. When a game arrives, each team member chooses whether to cooperate with others. In any given game, a team member prefers overall team cooperation to a lack of cooperation but individually prefers not to cooperate. Thus, the games have a prisoners’ dilemma-like structure.
This is not a simple repeated game: the stage games take on different forms. In some of the games, cooperation is beneficial to society (“good games”); in other games, cooperation is harmful (“bad games”). Socially optimal equilibria are those in which team members cooperate in good games only. Yet as Tirole (1988) conjectured, “…it may be worth tolerating some minor acts of detrimental co-operation to allow trust between employees to develop.” We show that the team-design problem is nontrivial because of interdependencies between the different stage games: equilibrium structure depends on the arrival frequencies of both good and bad games as well as their payoff structures.
Our analysis takes a design perspective. By adjusting team composition and assignments to various tasks, an organizational designer can control the incentives and behavior of the agents involved without observing actions in specific games or even whether a game occurs. This opens a new set of mechanism design problems, in which the dynamic set of tasks faced is controlled by the designer. This design perspective involves several dimensions.
One design dimension that we explore is team composition and durability. Some organizations, such as the military or police, keep teams together over time to foster cooperation. Others—e.g., diplomacy, retail, logistics, and services—frequently reshuffle members across teams or simply hire short-term contract workers for certain tasks, eliminating the repeated interaction necessary for cooperation. In which settings should teams be kept together versus reshuffled? Again, this issue is nontrivial as Tirole (1988) anticipated: “Fighting the detrimental cooperation by keeping the employees’ relationship short may also kill the beneficial co-operation.” We formalize this insight and identify conditions that make detrimental cooperation necessary to sustain socially beneficial cooperation.
@@
We also consider dynamic and reactive versions of the assignment of teams to tasks—taking full advantage of the dynamic aspect of dynamic design. Teams may be completing a certain task for some duration of time before being moved to another; they may also be moved in response to, say, an opportunity to cooperate in a situation where it is socially beneficial, e.g., after police officers find themselves in a dangerous spot. Dynamically and reactively reassigning teams to tasks further allows the organization to affect the random stream of opportunities to cooperate. We fully characterize the optimal structure of such assignments for a special case—when it is sufficiently easy to sustain cooperation in the bad game.
@@
Game theory, team theory, and mechanism design generally take the design problem as designing incentives for one particular (in some cases, dynamic) context. Here, we examine both the design of who faces which dynamic sequence of contexts, as well as with whom they play. [^3] There is a literature that examines how the diversity of teams affects their performance potential (e.g., see Page (2008)), but that is based on team heterogeneity, which is not considered here, and that literature does not consider any of the incentive issues or dynamics considered here. To our knowledge, this is the first paper to analyze the interaction of incentives across games, as well as the dynamic design of the sequence of games faced. This also provides a first analysis of treating the assignments of players to teams, and teams to games, as endogenous. This provides new angles for understanding team and organizational design.
@@
Our model is a stochastic game (introduced by Shapley, 1953; see Solan and Vieille, 2015 for a short review). Our model’s special structure—the players’ current actions do not affect the distribution over subsequent games—may be useful for extensions. This environment allows us to characterize equilibria, which are typically difficult to solve for (Solan and Vieille, 2015). Moreover, our analysis involves the design of a stochastic game, rather than analyzing a given one.
@@ 2 The Model
Player 1 C $(c,c)$ (a,c+d)(-a,c+d)
N (c+d,a)(c+d,-a) $(0,0)$
where $c,d>0$ and $a>c+d$. All players cooperating maximizes the total payoff, but a player can benefit by unilaterally deviating. The condition that $a>c+d$ ensures that the game satisfies Properties 5 (since then 0>c+da0>c+d-a) and 6 (since then $c+d-c\leq 0-(-a)$). The resulting cooperation benefit and deviation temptation are $c_{g}\equiv c$ and $d_{g}\equiv d$, respectively.
@@ 2.1 Cooperation
The different kinds of cooperation can be partially ordered by some social criterion. Let $V_{G}>0$ denote the social benefit from cooperation in game $G$, the good game, from the perspective of a social planner or designer. For example, this could be the social benefits to a society’s broader population of having a military or police force that cooperates when in dangerous situations, or the profits to a company from having a team cooperate on a productive task, etc. Similarly, $V_{B}<0$ denotes the social harm from cooperation in the bad game $B$. Thus, cooperation only in game $G$ is clearly most preferred; we call this optimal cooperation. Cooperation only in game $B$ is clearly the worst. The social value of total cooperation, $p_{G}V_{G}+p_{B}V_{B}$, and the social value of no cooperation, 0, lie in between the best and worst outcomes; their ordering depends on the relative social impact of cooperation in the two games and on their probabilities.
@@ 2.2 Player-Optimal Equilibria
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{B}/c_{B}$x\ xregion 2region 1region 3
@@
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{G}/c_{G}$x\ xregion 2region 1region 4
@@ 3 How Cooperation Depends on Team Durability
We model the reshuffling of teams (reducing team durability) as a probability $r$ with which the stochastic game of a team ends after each period. Given the continuum of workers, we assume that reshuffled players never face a past teammate and that there is no spreading of information and that players only condition on histories with their current team. [^7] This greatly simplifies the analysis; see, e.g., Ali and Miller (2016b) for a treatment where such contagion is possible. Our simplification also requires sufficiently many workers so that completely new teams can be formed and anonymity of workers outside their teams (deviations can be punished only by current team members). These conditions are more likely to hold—at least approximately—in large organizations, potentially leading to differences by organization size that we leave for future research. This can also be handled by hiring people part-time or by outsourcing and subcontracting tasks on short-term rotations. Consequently, a reshuffling rate $r$ is equivalent to changing the discount factor $\delta$ to $(1-r)\delta$; essentially, the designer can induce any effective discount factor of at most $\delta$ by varying team durability. Thus, formally, we characterize optimal reshuffling by examining the player optimal equilibria for a given team (with histories only applying to that team) as a function of discount factor $(1-r)\delta$.
@@
[…]
@@ 4 Assigning Teams to Tasks
A team’s history at time $\tau$ is now augmented with the history of assigned tasks, i.e., it is a vector $(t_{\tau^{\prime}},g_{\tau^{\prime}},\mathbf{a}_{\tau^{\prime}})]^{\tau-1}_{\tau^{\prime}=0}$ where $t_{\tau^{\prime}}$ is the task faced at time $\tau^{\prime}$. Let $\mathcal{H}_{\tau}$ denote all such histories and $\mathcal{H}\equiv\cup_{\tau}\mathcal{H}_{\tau}$. A player’s strategy indicates what the player who falls under a particular $\nu$ does as a function of the history, i.e., $\sigma:\mathcal{H}\times\{G,B\}\rightarrow\Delta(C,N)$, similar to before. Again, we track histories only within a team and consider equilibria that depend only on current team histories.
@@ 5.2.1 Good Games Are Observable
Because reactive assignments allow arbitrary (random) conditioning on full histories, the policy class is extremely rich. Our characterization pins down the global optimum within this high-dimensional space and shows that optimality is attained by a simple policy with only two building blocks (fully cooperative vs fully non-cooperative teams) and a minimal bad-block after good arrivals.
@@ 5.2.3 The Continuous Time Limit
We conclude our discussion on designer-optimal reactive team structures by considering how they depend on the length of a time period. In particular, we consider the continuous limit of our model; i.e., when time periods become infinitesimally small and, with them, the probability of a game occurring and the potential frequency of reassignments. We will see that whether the designer can improve upon static assignments as the period length shrinks depends on whether the arrival of good games is observable. This suggests that dynamic assignments in the unobservable case are only useful in situations where at most one good game arrives over a nontrivial span of time.
@@
We first note that as $T$ decreases, while the designer’s payoff from any fixed static team structure stays the same, the designer’s payoff from an optimal static team structure increases. The expected arrival rate of games is unchanged, but the timing of arrivals is more continuous rather than concentrated at the end of a full period. Under exponential discounting, a continuous stream of possible arrivals within, say, the next two periods is valued more highly than a single arrival exactly one period away. Thus, the same expected rate of future cooperation is effectively worth more. As a result, the designer can sustain cooperation with a smaller share of bad tasks, and the social value under the optimal static assignment increases as $T\to 0$.
Now consider reactive team structures and the case where the designer observes good games occurring (Section 5.2.1). The optimal reactive team structure does not change substantively: Consider a reactive assignment in which players always cooperate, which is part of an optimal reactive team structure. In the limit as $T\rightarrow 0$, following a good game, teams spend some fixed amount of time on the bad task before returning to the good task, where they wait until a good game arrives, as before. Furthermore, as was true in the $T=1$ case, the social value is higher when reactive assignments are possible compared to static assignments, as the designer can wait for the arrival of a good game. [^18] Thus, combining this conclusion with the observation in the previous paragraph, both the reactive and static assignments are improving, but the reactive is strictly better.
@@ 6 Concluding Remarks
7 Bells and Whistles
In closing, we discuss a few extensions.
Hybrid tasks.
When assigning teams to tasks, we focused on tasks that either had only good game or only bad games. Hybrid tasks that involve some random game arrivals could be accommodated, albeit with complications in notation and cases, but would yield similar results throughout. Also, in our model, the set of tasks is fixed. However, when supporting optimal cooperation is challenging, it can be advantageous to introduce new tasks that provide additional benefits from cooperation. These new tasks might not benefit society or directly increase organizational productivity, but they might help sustain cooperation in good games. For instance, scheduling “team-building” exercises, social events, and non-production related contests in which team members must cooperate with each other can help induce cooperation in other games as well. Many organizations use such team-building to enhance the loyalty that people feel towards the units in which they are involved, fostering cooperation. Although such events may be time-consuming and thus reduce the time that teams spend directly on productive activities, they may end up enhancing productivity and substitute for bad games by enabling cooperation in good games when that might have otherwise been infeasible.
Heterogeneity.
There may also be individual heterogeneity of team members in their preference to engage in collusive behavior. Adding individuals with a strong inclination to defect on bad games could curb collusion, but could also undermine productive cooperation. Similarly, “changing the culture” of an organization and increasing awareness of the negative consequences of collusion—that is, shifting team members’ preferences to be more averse to collusion on bad tasks—may erode the informal incentives that sustain cooperation on good tasks and productive teamwork.
Learning and coordination.
We have focused on the incentives for cooperation in our analysis. There can be other benefits from keeping teams together, such as players learning from each other and anticipating each other’s behaviors, etc. This provides additional reasons for keeping teams together, which could be incorporated into an extended model.
Monitoring.
Lastly, surveillance and other technologies also play an important role and may be modeled explicitly. Improved monitoring can make cooperation directly enforceable in good games and punishable in bad ones, thereby making incentive issues less relevant. Thus, organizations can benefit from investing in technologies to monitor behavior. For example, if cooperation in the good game can be directly enforced by a monitoring technology, reshuffling teams to reduce residual collusion makes good sense. However, Section 4 shows that an organization can benefit from the games’ strategic interdependencies. For example, consider an organization that prefers total cooperation to no cooperation. If no cooperation in the bad game is directly enforced via some new technology, then cooperation in the good game may be impaired, as players can now only punish deviation from cooperation via that game and not the bad game. All cooperation may thereby break down.
@@ Appendix B Proofs
Consider the full-specialization team structure $\lambda$ with $\lambda(\nu^{g}(t_{g}))=q_{g}$ for $t\in\{G,B\}$, $r_{G}=0$, and $r_{B}=1$, and first suppose $d_{G}\leq\frac{\delta}{1-\delta}a_{G}c_{G}$. Consider a team assigned to $\nu^{G}$. Property 2 implies that on-path aggregate stage-game payoffs are maximized when every player chooses grim-trigger strategy ‘Always cooperate unless any player did not cooperate in the past, at which point always do not cooperate.’ Furthermore, the resulting strategy profile constitutes an equilibrium: Facing the good game, the only game players of such team face, no player benefits from deviating since not cooperating leads to an increase in the stage-game payoff of $d_{G}$ and a decrease in future payoffs of $\frac{\delta}{1-\delta}a_{G}c_{G}$ and by assumption and since not cooperating if someone deviated in the past is a Nash equilibrium (an implication of Property 3), i.e., the punishment stage is a Nash equilibrium. Consider now a team assigned to $\nu^{B}$. As $r_{B}=1$, players’ discount factor is 0 and so they play the unique stage game equilibrium, which is no cooperation. Thus, there is cooperation only on good games.
@@
Consider a class of reactive assignments $m$ parameterized by $w_{gg^{\prime}}\in[0,1]$ for $g,g^{\prime}\in\{G,B\}$ defined as follows. Consider any state $(t_{\tau},g_{\tau})]_{\tau=0}^{K}$ with current task $t_{K}=t_{g}$; then let $m([(t_{\tau},g_{\tau})]_{\tau=0}^{K})(t_{g}^{\prime})=w_{gg^{\prime}}$ for $w_{gg^{\prime}}$ such that $w_{gG}+w_{gB}=1$ for all $g\in\{G,B\}$. Note that for $w_{Gg}=w_{Bg}=\nu^{coop}(t_{g})$, then $m$ corresponds to the same static assignment as $\nu^{coop}$.
@@
Let $\overline{g},\underline{g}$ be such that $d_{\overline{g}}>d_{\underline{g}}$, i.e., it is harder to incentivize cooperation in game $\overline{g}$. The conditions of the proposition imply $a_{G}c_{G}<a_{B}c_{B}$. Therefore, in an assignment $m$ that is not the same static assignment as $\nu^{coop}$ the average benefit from future cooperation can be made strictly larger when teams are in game $\overline{g}$ than when they are in game $\underline{g}$ by sending them more frequently to the bad task when in $\overline{g}$, i.e., by increasing $w_{\overline{g}B}$. Furthermore, this can be done while maintaining the same unconditional benefit from future cooperation in this modified $m$ and original $\nu^{coop}$ by decreasing $w_{\underline{g}B}$ according to Equation 8. By continuity and if $d_{\overline{g}}$ is strictly less than $d_{\underline{g}}$, there are thus values of $w_{BG},w_{GB}$ (with $w_{BB}=1-w_{BG},w_{GG}=1-w_{GB}$) such that the benefit from future cooperation is strictly greater than the deviation payoff in both tasks, which allows the designer to strictly improve the social value, as in the proof of Lemma 3. ∎